A person with a near-point distance of 32 cm is able to get a usable angular magnification of 12 with a particular magnifying glass. Her friend is only able to get a usable angular magnification of 7. What is the friend’s near-point distance?
Answer #1

we know that, the Angular magnification of the eyepiece

m = n/f where n is the near point and f is the focal length

f =n/m = 32/12 =2.66666667

again from the same formula we have, n = f x m for the friend, n2 = f x m2

=  2.66666667*7 =  18.6666667 cm  

the friend’s near-point distance = 18.6666667 cm answer